# The Cone Restriction: An old Approach Revisited

### v4

@@ 1 Introduction
$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p},\frac{1}{p}+\frac{1}{p^{\prime}}=\frac{1}{q}+\frac{1}{q^{\prime}}=1$ (1.2)
@@
where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$.
@@
[…]
By a short duality argument, we see that conjecture 1.1 is equivalent to
@@
Let $n\geq 3$. For all $p,q>0$ satisfy (1.2), there holds
$\|Ef\|_{p}\lesssim_{n,p,q}\|f\|_{q},\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$ (1.3)
@@
For $n\geq 3$, (1.3) holds whenever
$q\geq p>\left\{\begin{array}[]{rcl}4&\mbox{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mbox{if}&n>3\quad odd\\ 2\frac{3n}{3n-4}&\mbox{if}&n>3\quad even\end{array}\right.$ (1.4)
@@
For $n\geq 3$, (1.3) holds whenever the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mbox{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mbox{if}&k\geq 3\end{array}\right.$
@@
By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [6]), we see that
@@
For $n\geq 3$, (1.3) holds whenever $q\geq p$ and (1.6) holds.
Guth [8, 9] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids. Later, his results were subsequently refined by Wang[10], Wang-Wu[11], Hickman–Rogers [12], Hickman–Zahl [13], and Guo-Wang-Zhang [14]. In [12, 13] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [15, 16, 17], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [18, 19]), the Schrödinger maximal estimate (see [20]), and cone restriction (see [4]).
In this article, we refine the results presented in [4] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [12, 13]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [4] as a recursive process. Compared to the paraboloid case addressed in [12, 13], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its $(n-1)$- dimensional ancestor. We will provide a more detailed discussion on this topic in Section 3.
@@ 2.1 Notational Conventions
$\delta:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We say $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$. Let #$\#A$ denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}$ denote an arbitrary ball in $\mathbb{R}^{n}$ of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$. We say $A^{\prime}\subseteq A$ is a refinement of $A$ if #$\#A\lessapprox\#A^{\prime}$. We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$. We say $f$ is essentially supported in $\Omega$ if $f(x)\lesssim$ RapDec($R$) $\|f\|_{2}$. We say $Ef$ is essentially supported in $\Omega$ if $Ef(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
@@ 2.2 Wave Packet Decomposition
The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from section 2.2 and 5.2.1 of [4].
@@
Like the case of paraboloid [12], there is also some local versions of the $L^{2}$-orthogonality for cone, which are a direct corollary of lemma 2.1:
@@
One can verify that $Eg\left(x\right)=E\tilde{g}\left(\tilde{x}\right)$ where $x=y+\tilde{x}$ for some $\tilde{x}\in B\left(0,\rho\right)$. Next, we break $\tilde{g}$ into scale $\rho$ wave packets
@@
Let $T^{L}_{\zeta,w}\left(y\right)$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$). Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
@@
Since each $\left(\zeta,w,L\right)$ belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
@@
The degree of the transverse complete intersection is defined as ${\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\ }$.
@@
In secton 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
@@ 2.3 Some Kakeya-type geometric results
$N_{r^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{r^{1/2+{\delta}_{t}}_{t}}Z_{t}$
$r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter $\left(t-s+1\right)$ is referred to as the level of multigrain. The degree of $\overrightarrow{S}$ is defined as ${\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\ }$. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$. Finally, we write
@@
for any $L\in\mathcal{L}$, then what is the maximum possible value of $\#\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
@@
#$\#\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
@@
$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]$
For $m<i\leq n-1$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]$
@@
$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
One can observe the second condition in definition 2.15 is weaker than lemma 2.5. However, this is good enough for our arguments. In [13, 19, 25] a non-trivial extension of Theorem 2.14, which is a corollary to lemma 2.11 in [17], was used to establish a new result on the Fourier restriction, the Bochner-Riesz problem, and the Stein’s square functions respectively. By setting $r_{0}=r_{1}$ in this extension (see [13], Lemma 3.7), one has
Lemma 2.16 (Nested polynomial Wolff axioms).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d,\ \ \forall i\in\mathbb{N}\cap\left[m,n\right)$
@@
#$\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
In [13, 19, 25] $S_{m}$ was passed back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus
$\mathrm{\Theta}\left[{\overrightarrow{S}}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
Since $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-1\right)/2}$ in the case of paraboloids, this bound gives a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$. However, in the case of cone this doesn’t guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$. This is the reason for passing $S_{m}$ back to $S_{n-1}$ rather than $S_{n}$ in our arguments.
@@ 2.4 $k$-broad Norms
Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about .${\left\|.\right\|}_{BL^{p}_{k,A}}$.
@@
Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If ${\mathrm{deg}\boldsymbol{Z}\lesssim 1\ }$ and ${dim\boldsymbol{Z}\ }$ $\leq k-1$, then
@@
If $\boldsymbol{\mathrm{Z}}$ is a plane, then the existence of such a decomposition follows from the Lemma 2.6. However, this is not easy for ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular. A rigorous version of (2.4) and (2.5) is given by the following statement. See [9, 4, 7] for details.
@@
$\ \ \sum_{b\in\mathcal{B}}{{\left\|g^{\prime}_{b}\right\|}^{2}_{2}}\lesssim{\left\|g^{\prime}\right\|}^{2}_{2}$
@@
[…]
@@
To obtain the best possible estimate in the theorem 1.5, $v$ should satisfy

### v5

@@ 1 Introduction
$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p}$ (1.2)
@@
where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$, and $p^{\prime},q^{\prime}$ are the Hölder conjugates of $p,q$.
@@
$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}|\xi|)}f(\xi)\,d\xi\$
where $B^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that conjecture 1.1 is equivalent to
@@
Let $n\geq 3$. For all $p,q>0$ satisfy (1.2),
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
@@
Let $n\geq 3$. For all $p,q>2$ satisfy
$q\geq p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$.
there holds
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
@@
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
then
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
@@
By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [17]), one has
@@
Let $n\geq 3$. For all $q\geq p>2$ satisfy (1.4), there holds
$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids. Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7]. In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, HRZ, 24], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [21, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
In this article, we refine the results presented in [15] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [11, 12]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [15] as a recursive process. Compared to the paraboloid case addressed in [11, 12], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor. We will provide a more detailed discussion on this topic in Section 3.
@@ 2.1 Notational Conventions
$\delta_{0}:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We say $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$. Let #A\#A denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}$ denote an arbitrary ball in $\mathbb{R}^{n}$ of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$. We say $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$. We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$. We say $f$ is essentially supported in $\Omega$ if $f(x)\lesssim$ RapDec($R$) $\|f\|_{2}$. We say $Ef$ is essentially supported in $\Omega$ if $Ef(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
@@ 2.2 Wave Packet Decomposition
The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
@@
Like in the case of paraboloid [11], there are also some local versions of the $L^{2}$-orthogonality for the cone, which are direct corollaries of Lemma 2.1:
@@
[…]
@@
Let $T^{L}_{\zeta,w}\left(y\right)\$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$). Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
@@
Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
@@
The degree of the transverse complete intersection is defined as $\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\$}.
@@
In Section 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
@@ 2.3 Some Kakeya-type geometric results
$B(y_{i},r_{i})\cap B(y_{i+1},r_{i+1})\neq\emptyset,\quad\forall i$
$N_{Cr^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{Cr^{1/2+{\delta}_{t}}_{t}}Z_{t}\mathrm{\ for\ some\ }C>0$
$0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter $\left(t-s+1\right)$ is called the multigrain level. The degree of $\overrightarrow{S}$ is defined as $\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\$}. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$. Finally, we write
@@
for any $L\in\mathcal{L}$, then what is the maximum possible value of $\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
@@
$\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
@@
$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
For $m<i\leq n$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]\$
@@
$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\ \mathrm{\Theta}\left[{\overrightarrow{S}}^{n-1}_{m}\right]$
One can observe that the second condition in Definition 2.15 is weaker than Lemma 2.5. However, this is good enough for our arguments. In [12, 8, 10] a nontrivial extension of Theorem 2.14, which is a corollary to Lemma 2.11 in [24], was used respectively to establish a new result on the Bochner-Riesz problem, Stein’s square function, and the Fourier restriction for paraboloid. The result is shown below:
Lemma 2.16 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
@@
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n}\left(\prod^{n-1}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-1}{2}}_{n}$
A direct corollary of this lemma is
Corollary 2.17.
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
Proof.
Apply Lemma 2.16 to ${\overrightarrow{S}}:=(S_{m},...,S_{n-1},({\mathbb{R}}^{n},B(y_{n-1},r_{n-1})))$. ∎
In [12, 8, 10] $S_{m}$ is traced back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus Lemma 2.16 gives
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
Since the tubes have $\sim R^{\left(n-1\right)/2}$ directions in the case of paraboloids, this bound yields a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$. However, in the case of cone, this does not guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$. This is the why we need Corllary 2.17
@@ 2.4 $k$-broad Norms
Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about ${\left\|.\right\|}_{BL^{p}_{k,A}}$.
@@
Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If $\mathrm{deg}\boldsymbol{Z}\lesssim 1\$} and $dim\boldsymbol{Z}\$} $\leq k-1$, then
@@
If $\boldsymbol{\mathrm{Z}}$ is a plane, then such a decomposition exists due to Lemma 2.6. However, this question will be tricky if ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular. A rigorous version of (2.4) and (2.5) is given by the following statement. See [6, 15, 4] for details.
@@
[…]
@@
$p\left(k,n\right)<p\leq\frac{2n}{n-2},p\left(k,n\right):=\left\{\begin{array}[]{c}2+\frac{2}{n-2},\ if\ 2\leq k\leq 3\\ 2+\frac{4}{2n-k-1},\ if\ k>3\end{array}\right$. (2.8)
@@
To obtain the best possible estimate in Theorem 1.5, $v$ should satisfy